# FedHK-MVFC: Federated Heat Kernel Multi-View Clustering

### v1

@@ Abstract
In the realm of distributed AI and privacy-focused medical applications, we propose a framework for multi-view clustering that links quantum field theory with federated healthcare analytics. Our method uses heat-kernel coefficients from spectral analysis to convert Euclidean distances into geometry-aware similarity measures, capturing the structure of diverse medical data. We lay this out through the Heat Kernel Distance (HKD) transformation with convergence guarantees. Two algorithms are developed: Heat Kernel-Enhanced Multi-View Fuzzy Clustering (HK-MVFC) for central analysis, and Federated Heat Kernel Multi-View Fuzzy Clustering (FedHK-MVFC) for secure, privacy-preserving learning across hospitals using differential privacy and secure aggregation to facilitate HIPAA-compliant collaboration. Tests on synthetic datasets of cardiovascular patients show an 8-12% increase in clustering accuracy, 70% reduced communication, and 98.2% efficiency retention over centralized methods. Validated on 10,000 patient records across two hospitals, it proves useful for collaborative phenotyping involving ECG, cardiac imaging, and behavioral data. Our theoretical contributions include update rules with proven convergence, adaptive view weighting, and privacy-preserving protocols. This presents a new standard for geometry-aware federated learning in healthcare, turning advanced math into workable solutions for analyzing sensitive medical data while ensuring both rigor and clinical relevance.
@@ 1 Introduction
The increasing prevalence of heterogeneous data sources has spurred significant interest in multi-view (MV) clustering, wherein a single entity is characterized by multiple distinct feature sets or “views.” In such contexts, traditional clustering methods often fall short, necessitating sophisticated approaches that can effectively integrate complementary information across diverse views while maintaining coherence in the resulting partition. In addition, the rapid evolution of large-scale artificial intelligence (AI) systems - from centralized language models to decentralized, agent-based architectures - has fundamentally shifted how learning is conceptualized and deployed. In these new paradigms, learning must occur not only across heterogeneous data views but across distributed agents, each operating within their own local context, resource constraints, and knowledge representation. This shift demands a theoretical and algorithmic framework that accommodates structural diversity, communication limitations, and the intrinsic geometry of information flow.
@@
In this work, we propose a novel soft clustering framework that replaces conventional Euclidean distance (CED) with kernelized geometric measures derived from H-KC – tools rooted in QFT and spectral analysis. By modeling similarity via heat diffusion over data manifolds, our approach provides a richer, geometry-aware mechanism for aggregating and aligning distributed views. Moreover, the theoretical underpinnings from QFT offer a natural interpretation of influence propagation – a notion increasingly relevant in the context of collaborative AI and agent-based systems.
Consider the case of multiple hospitals collaboratively monitoring cardiovascular health through FL – one through ECG signals, another via MRI imaging, others through wearables, EHRs, or genetic data. While CED treats each view as flat and uniform, it fails to capture the curved anatomical structures in imaging or the relational dependencies in medical histories. The HKD, in contrast, models how “information” or similarity diffuses across each hospital’s local data manifold—honoring intrinsic structures. This enables the clustering of patients who may appear dissimilar numerically but share underlying physiological patterns across modalities.
@@
Our method enables flexible, privacy-preserving learning across multiple views and agents, with theoretical guarantees and practical scalability. It bridges foundational mathematics with contemporary machine learning (ML) architecture, contributing a unified lens to interpret and design decentralized intelligent systems (DIS). In summary, the contributions of this study include:
@@
    In the era of modular AI, where not only data views but distributed agents must harmonize across diverse sources, we propose a unified theoretical framework bridging MV learning, kernel geometry, and quantum field theory (QFT) to support scalable, interpretable federated learning (FL).
@@
    Just as large language models (LLMs) or “LLM-agents” encode semantic geometry in high-dimensional manifolds, this work proposes heat kernel-based soft clustering to capture local geometric structures across decentralized views. In other words, this paper revisits the principles of information diffusion, not only across data views but across distributed agents-drawing connections from classical field theory to the most recent breakthroughs in neural language model.
@@
    We introduce a novel centralized clustering algorithms that bridge the Heat Kernel-Enhanced Multi-View Fuzzy Clustering (HK-MVFC).
  4. 4.
    The proposed HK-MVFC algorithm employs H-KC to transform the CED metric into an HKD, enabling more effective capturing of complex relationship between data points and cluster centers across multiple views. Here, HK-MVFC incorporates view-specific weighting factors to balance the contribution of each view and treats feature-specific weights across views are equally important.
  5. 5.
    We implement the mathematical formulation for federated HK-MVFC, abbreviated as FedHK-MVFC. The federated variant of HK-MVFC employs a parameter synchronization protocol that preserves data privacy.
@@ 2 Literature Review
In this section, we review the relevant literature on federated learning, clustering, and their intersections.
Federated learning (FL) has emerged as a paradigm for collaborative model training across decentralized data sources while preserving privacy [1, 2]. Early works focused on federated averaging (FedAvg) and its variants, addressing communication efficiency and robustness to data heterogeneity. Recent advances include personalization strategies [3, 4], secure aggregation [5], and applications in healthcare, finance, and IoT.
Clustering algorithms, such as k-means, hierarchical clustering, and fuzzy c-means (FCM) [6], are foundational in unsupervised learning. Multi-view clustering extends these methods to data with multiple feature representations, leveraging complementary information for improved accuracy [7, 8]. Kernel-based approaches, including spectral clustering and kernel k-means, address non-linear relationships and manifold structures [9, 10].
Distributed clustering integrates clustering algorithms with distributed computing frameworks to scale to large datasets and decentralized environments [11, 12]. Federated clustering combines FL principles with clustering, enabling privacy-preserving unsupervised learning across clients [13, 14]. Existing federated clustering methods often rely on communication-efficient protocols and local model updates, but face challenges in handling heterogeneous data distributions and aligning cluster assignments across clients.
Recent works have explored kernel methods and manifold learning in federated settings [15, 16], highlighting the importance of geometry-aware similarity measures. However, most approaches use conventional distance metrics and lack theoretical guarantees for convergence and consistency in multi-view, distributed contexts.
Our work builds on these foundations by introducing heat-kernel coefficients and kernelized distances for multi-view federated clustering, bridging quantum field theory concepts with practical distributed learning algorithms.
2.1 Federated Learning
Federated learning (FL) enables decentralized model training by aggregating updates from multiple clients, each operating on local data. The core mechanism involves iterative communication between clients and a central server, where clients compute local model updates and the server aggregates these updates to refine the global model. This approach mitigates privacy risks associated with centralized data storage and is particularly suited for sensitive domains such as healthcare, finance, and mobile applications.
Key challenges in FL include handling data heterogeneity, ensuring communication efficiency, and maintaining model convergence. Techniques such as federated averaging (FedAvg) [1], secure aggregation [5], and personalization strategies [3, 17] have been developed to address these issues. Recent research also explores robust aggregation methods, differential privacy, and federated optimization algorithms to further enhance FL’s scalability and reliability.
In the context of clustering, federated learning introduces additional complexity due to the unsupervised nature of the task and the need to align cluster assignments across clients. Existing federated clustering approaches often rely on communication-efficient protocols and local model updates, but face challenges in managing heterogeneous data distributions and achieving consistent clustering results. Our work builds on these foundations by integrating geometry-aware kernel methods and heat-kernel coefficients to improve clustering performance in federated multi-view settings.
2.2 Clustering
Clustering algorithms are central to unsupervised learning, aiming to partition data into groups based on similarity. The classical k-means algorithm assigns each data point to the nearest cluster centroid, minimizing intra-cluster variance. Hierarchical clustering builds nested clusters by either agglomerative (bottom-up) or divisive (top-down) strategies, producing a dendrogram that reveals data structure at multiple scales. Fuzzy c-means (FCM) [6] generalizes k-means by allowing each data point to belong to multiple clusters with varying degrees of membership, controlled by a fuzzifier parameter.
Multi-view clustering extends these paradigms to datasets where each sample is described by multiple feature sets (views). By integrating information from different views, multi-view clustering can achieve more robust and accurate partitions [7, 18]. Approaches include co-regularization, consensus clustering, and weighted integration, each addressing the challenge of view heterogeneity.
Kernel-based clustering methods, such as spectral clustering and kernel k-means, leverage kernel functions to capture non-linear relationships and manifold structures in data [9, 10]. Spectral clustering uses the eigenvectors of similarity matrices to embed data into a lower-dimensional space before applying conventional clustering. Kernel k-means replaces the Euclidean distance with kernel-induced similarity, enabling clustering in high-dimensional feature spaces.
Recent advances focus on adaptive weighting of views, automatic selection of kernel parameters, and scalable algorithms for large multi-view datasets. Despite progress, challenges remain in handling noisy views, aligning cluster assignments across views, and providing theoretical guarantees for convergence and consistency.
2.3 Distributed Clustering
Distributed clustering refers to clustering algorithms designed to operate on data that is partitioned across multiple nodes or clients, often in large-scale or privacy-sensitive environments. In such settings, data cannot be centralized due to privacy, bandwidth, or computational constraints. Distributed clustering methods typically fall into two categories: (1) centralized coordination, where a server aggregates local clustering results or model parameters from clients, and (2) decentralized protocols, where clients communicate directly or via peer-to-peer networks to reach consensus.
Federated clustering is a specialized form of distributed clustering that emphasizes privacy preservation and minimal data sharing. In federated clustering, each client computes local cluster assignments or updates based on its own data, and only shares model parameters (such as cluster centroids or membership matrices) with a central server. The server aggregates these updates—using techniques like weighted averaging or secure aggregation—to refine a global clustering model, which is then redistributed to clients for further local updates. Communication-efficient protocols, such as periodic synchronization or compressed updates, are often employed to reduce bandwidth usage and improve scalability.
Key challenges in distributed and federated clustering include handling heterogeneous data distributions, aligning cluster assignments across clients, ensuring convergence, and maintaining privacy. Recent advances leverage kernel methods, consensus algorithms, and adaptive weighting to address these issues, enabling robust clustering in decentralized environments without compromising data privacy.
The evolution of these techniques has led to more sophisticated frameworks that can dynamically adapt to varying data characteristics and client capabilities. For detailed discussions on distributed clustering, refer to [11, 12].

### v2

@@ Abstract
In the realm of distributed artificial intelligence (AI) and privacy-focused medical applications, this paper proposes a multi-view clustering framework that links quantum field theory with federated healthcare analytics. The method uses heat kernel coefficients from spectral analysis to convert Euclidean distances into geometry-aware similarity measures that capture the structure of diverse medical data. The framework is presented through the heat kernel distance (HKD) transformation, which has convergence guarantees. Two algorithms have been developed: The first, Heat Kernel-Enhanced Multi-View Fuzzy Clustering (HK-MVFC), is used for central analysis. The second, Federated Heat Kernel Multi-View Fuzzy Clustering (FedHK-MVFC), is used for secure, privacy-preserving learning across hospitals. FedHK-MVFC uses differential privacy and secure aggregation to enable HIPAA-compliant collaboration. Tests on synthetic cardiovascular patient datasets demonstrate increased clustering accuracy, reduced communication, and retained efficiency compared to centralized methods. After being validated on 10,000 synthetic patient records across two hospitals, the methods proved useful for collaborative phenotyping involving electrocardiogram (ECG) data, cardiac imaging data, and behavioral data. The proposed methods’ theoretical contributions include update rules with proven convergence, adaptive view weighting, and privacy-preserving protocols. These contributions establish a new standard for geometry-aware federated learning in healthcare, translating advanced mathematics into practical solutions for analyzing sensitive medical data while ensuring rigor and clinical relevance.
@@ 1 Introduction
The proliferation of heterogeneous data sources has catalyzed a surge of interest in multi-view (MV) clustering, a methodology in which a single entity is delineated by multiple distinct feature sets or ”views.” In such contexts, conventional clustering methods frequently prove inadequate, underscoring the necessity for advanced approaches that can seamlessly integrate complementary information across diverse perspectives while preserving coherence in the resulting partition. Furthermore, the rapid evolution of large-scale artificial intelligence (AI) systems—from centralized language models to decentralized, agent-based architectures—has fundamentally shifted the conceptualization and deployment of learning [montes2019distributed]. In these novel frameworks, learning must occur not only across heterogeneous data views but also across distributed agents, with each agent operating within its own local context, resource constraints, and knowledge representation. This paradigm shift necessitates a theoretical and algorithmic framework that can accommodate structural diversity, communication limitations, and the intrinsic geometry of information flow.
In the context of prospective Internet applications, such as healthcare monitoring, smart cities, and distributed sensor networks, data is frequently collected and processed across a multitude of decentralized nodes [kremenova2019decentralized]. The utilization of conventional clustering methods is hindered by the inherent characteristics and structures of the numerous sources of data. These data sources exhibit varied characteristics and structures, which hinder the efficacy of traditional clustering methods that rely on flat Euclidean representations. This challenge has been thoroughly documented in the research literature [wang2004improving, jing2007entropy, cleuziou2009cofkm, fang2023comprehensive]. Healthcare serves as a prime example of this phenomenon. Patient data may be distributed across various hospitals, each collecting different types of information (e.g., imaging, physiological signals, and electronic health records) [fang2024decentralised]. Conventional clustering methodologies that employ conventional Euclidean distance (CED) may encounter challenges in accurately capturing the intricate relationships and inherent geometries present in such heterogeneous datasets. Consequently, there is an urgent need for clustering algorithms that can effectively integrate and analyze multi-view data in a decentralized manner while respecting the privacy and autonomy of each data source.
Federated learning (FL) has emerged as a promising paradigm for collaborative model training across decentralized data sources while preserving privacy [pei2024review]. In the state of Florida, a consortium of clients (e.g., hospitals) collaboratively trains a global model without sharing their raw data, thus addressing privacy concerns and regulatory constraints [abbas2024federated]. However, extant FL approaches frequently depend on conventional distance metrics and are deficient in their inability to capture the intrinsic geometries of multi-view data. This limitation impedes their efficacy in scenarios where data is heterogeneous and distributed across multiple agents.
1.1 Research Problem and Objectives
The fundamental research problem addressed in this work concerns the development of a unified framework for privacy-preserving multi-view clustering that simultaneously satisfies three critical requirements: (1) geometric awareness—capturing the intrinsic manifold structure of heterogeneous data that conventional Euclidean distance metrics fail to represent, (2) federated scalability—enabling distributed learning across autonomous agents without centralized data aggregation, and (3) theoretical rigor—providing mathematically proven convergence guarantees and privacy preservation mechanisms. Existing multi-view clustering approaches predominantly rely on Euclidean distance, which assumes linear relationships and uniform feature scaling, thereby failing to capture the complex geometric structures inherent in real-world heterogeneous datasets. Furthermore, federated learning frameworks for multi-view data lack principled mechanisms for aggregating complementary information across distributed views while maintaining privacy and communication efficiency.
To address these challenges, this research pursues the following specific objectives: First, we develop a heat kernel-based geometric measure that transforms conventional distance computations into geometry-aware similarity assessments, enabling effective clustering on non-linear manifolds and varying cluster densities. Second, we formulate a federated multi-view clustering algorithm that extends the centralized heat kernel framework to distributed settings, incorporating differential privacy mechanisms and secure aggregation protocols tailored for clustered network architectures. Third, we establish theoretical foundations including convergence guarantees, privacy bounds, and communication complexity analysis for the proposed federated clustering framework. Fourth, we validate the framework through comprehensive experimental evaluation on synthetic multi-view datasets representative of network-generated data, demonstrating substantial improvements over state-of-the-art baselines in clustering accuracy, communication efficiency, and privacy preservation.
@@ 1.1 Research Problem and Objectives
In this work, we propose a novel soft clustering framework that replaces conventional Euclidean distance (CED) with kernelized geometric measures derived from H-KC. These tools are rooted in QFT and spectral analysis. By modeling similarity via heat diffusion over data manifolds, our approach provides a richer, geometry-aware mechanism for aggregating and aligning distributed views. Furthermore, the theoretical underpinnings from QFT offer a natural interpretation of influence propagation, a notion that is increasingly relevant in the context of collaborative AI and agent-based systems.
@@
Our method enables flexible, privacy-preserving learning across multiple views and agents, with theoretical guaranties and practical scalability. It bridges foundational mathematics with contemporary machine learning (ML) architecture, contributing a unified lens to interpret and design decentralized intelligent systems (DIS). In summary, the contributions of this study include:
@@
    In the contemporary era of modular artificial intelligence (AI), where the harmonization of not only data views but also distributed agents across diverse sources is paramount, we propose a unified theoretical framework that integrates maximum-volume (MV) learning, kernel geometry, and quantum field theory (QFT) to facilitate scalable, interpretable federated learning (FL). We introduce the concept of heat-kernel coefficients (H-KC) as a novel mechanism to capture local geometric structures across decentralized views and agents, thereby enabling effective information diffusion in DIS. This framework provides a principled basis for designing algorithms that respect both data heterogeneity and agent autonomy.
@@
    The present study proposes a methodology for the analysis of local geometric structures across decentralized views. This methodology is based on heat kernel-based soft clustering, a technique that has been employed in the study of large language models (LLM) and ”LLM-agents,” which encode semantic geometry in high-dimensional manifolds. The present study proposes an extension to the concept of centralized clustering algorithms, with the objective of establishing a framework to bridge the Heat Kernel-Enhanced Multi-View Fuzzy Clustering (HK-MVFC). The mathematical formulation for federated HK-MVFC, abbreviated as FedHK-MVFC, is implemented. This algorithm facilitates the collaborative clustering of multi-view data by multiple clients (e.g., hospitals) while preserving privacy and minimizing communication overhead. The theoretical analysis provided herein demonstrates convergence properties and robustness to data heterogeneity.
@@
    As medical datasets of this nature are not publicly available, we have generated synthetic datasets that mimic the characteristics of real-world multi-view medical data. We contribute a novel synthetic dataset with non-linear relationships across views, designed to evaluate the effectiveness of geometry-aware clustering methods.
The remainder of this paper is organized as follows: In Section 2, an examination of extant literature pertaining to federated learning, multi-view clustering, and heat kernel methods is conducted. The subsequent sections (3-4) delineate the mathematical formulation of the proposed HK-MVFC and FedHK-MVFC algorithms. As delineated in Section 5, the proposed Fed-HKMVFC concept is predicated on a privacy-preserving framework. In Section 6, experimental results are presented on synthetic datasets. These results are compared with baseline methods. Conclusively, Section 7 culminates in a discourse on the ramifications and prospective avenues for further research.
@@ 2 Literature Review
Federated learning (FL) has emerged as a paradigm for collaborative model training across decentralized data sources while preserving privacy [mcmahan2017communication, konevcny2016federated]. Early works focused on federated averaging (FedAvg) and its variants to address communication efficiency and robustness to data heterogeneity. Recent advances include personalization strategies [smith2017federated, li2021ditto], secure aggregation [bonawitz2017practical], and applications in healthcare, finance, and the Internet of Things (IoT). Clustering algorithms, including k-means, hierarchical clustering, and fuzzy c-means (FCM) (see reference [bezdek2013pattern]), are foundational in unsupervised learning. Multi-view clustering extends these methods to leverage complementary information across multiple feature representations (see references [xu2013survey, chen2020multi]). Kernel-based approaches, such as spectral clustering and kernel k-means, address non-linear relationships and manifold structures (see references [ng2001spectral, dhillon2004kernel]). Distributed clustering integrates clustering algorithms with distributed computing frameworks to scale to large datasets and decentralized environments [guha2003clustering, xu2015comprehensive]. In contrast, federated clustering combines FL principles with clustering to enable privacy-preserving unsupervised learning across clients [brisimi2018federated, duan2021fedgroup]. However, existing methods often rely on conventional distance metrics and face challenges in handling heterogeneous data distributions and aligning cluster assignments across clients. Recent works have explored kernel methods and manifold learning in federated settings (see [chen2022federated, liu2023federated]), highlighting the importance of geometry-aware similarity measures. However, most approaches lack theoretical guaranties and practical scalability in complex multi-view scenarios.
In centralized multi-view clustering, Yang and Sinaga [yang2021collaborative] introduced Co-FW-MVFCM (Collaborative Feature-Weighted Multi-View Fuzzy C-Means), which enables collaborative learning across data views through adaptive feature weighting. While Co-FW-MVFCM demonstrates effectiveness in centralized environments, it lacks mechanisms for distributed deployment and cannot handle geometrically complex data structures such as non-convex clusters or manifold-embedded patterns. Similarly, Zhou et al. [zhou2025mvwecm] proposed MvWEC (Multi-view Weighted Evidential C-Means), incorporating evidential reasoning to manage uncertainty in multi-view clustering. Despite its theoretical contributions, MvWEC still relies on traditional distance measures that may not fully exploit the geometric structure of multi-view data, particularly when views exhibit complementary rather than redundant information.
The extension to federated settings introduces additional complications. Yang et al. [yang2024federated] proposed federated k-means clustering based on feature weighting, addressing privacy preservation and communication efficiency. However, this approach inherits the fundamental limitation of k-means—dependence on Euclidean distance—which restricts its capability to capture complex relationships in multi-view data. Liu et al. [liu2023federated] developed federated probabilistic preference learning for privacy-preserving multi-domain recommendation based on co-clustering. While addressing privacy concerns through local computation and secure aggregation, existing co-clustering methods often depend on conventional distance metrics, struggle with data heterogeneity across clients, and lead to inconsistent clustering results.
Recent tensorized approaches have extended single-view clustering to multi-view settings. Liu et al. [liu2024adaptively] introduced adaptively weighted multi-view tensorized clustering, leveraging tensor decomposition for view integration. However, the underlying distance computations remain Euclidean-based, limiting effectiveness on datasets with intrinsic geometric complexity. Furthermore, most tensor-based methods [chen2011tw, xu2016weighted, zhang2018tw, sinaga2024rectified, kaushal2024weighted] do not consider distributed or federated learning settings, restricting their applicability in privacy-sensitive scenarios such as healthcare and finance.
Despite significant advancements in federated learning and multi-view clustering, several fundamental limitations persist when integrating heterogeneous data across distributed environments. The principal challenge stems from the conventional distance metrics employed by existing methods, which fail to capture the intrinsic geometric structures inherent in multi-view data. Traditional approaches predominantly rely on Euclidean distance, which assumes linear relationships and uniform feature scaling—assumptions that are frequently violated in real-world heterogeneous datasets. This metric inadequacy becomes particularly problematic when dealing with non-linear manifolds, varying cluster densities, and complex topological structures common in medical imaging, social networks, and genomic data.
In healthcare applications, these limitations manifest critically. Multi-institutional patient data often exhibit complex geometric patterns across modalities (ECG signals, imaging, genomic profiles), heterogeneous statistical distributions due to demographic variations, and strict privacy constraints preventing raw data sharing. Existing methods fail to simultaneously address these requirements: geometry-aware similarity measures for capturing intrinsic data structures, federated learning protocols for privacy-preserving collaboration, and robust aggregation mechanisms for handling heterogeneous client distributions.
2.1 Research Gaps and Motivation
The comprehensive analysis of existing literature reveals three critical research gaps that motivate the present work:
Gap 1: Geometric Structure Ignorance. Current multi-view clustering methods, both centralized and federated, predominantly employ Euclidean distance metrics that fail to capture the intrinsic manifold geometry of complex data distributions. This limitation is particularly severe when data resides on non-linear manifolds or exhibits varying local densities—scenarios common in medical imaging, genomic analysis, and sensor network data. While kernel methods have been explored in single-view settings, their principled integration into multi-view federated frameworks remains unexplored, especially with theoretical foundations grounded in quantum field theory and spectral graph analysis.
Gap 2: Multi-View Heterogeneity in Federated Settings. Existing federated clustering approaches lack robust mechanisms for handling heterogeneous multi-view data across distributed clients. Current methods either aggregate views naively without considering their complementary nature or employ view-specific models that fail to leverage cross-view consistency. The challenge intensifies when different clients possess varying numbers of views with distinct statistical properties, a scenario frequently encountered in multi-institutional healthcare collaborations where data collection protocols differ across sites.
Gap 3: Privacy-Performance Trade-off. While differential privacy and secure aggregation techniques exist for federated learning, their integration with geometry-aware multi-view clustering introduces unique challenges. Existing privacy-preserving methods either sacrifice clustering accuracy significantly or require prohibitive communication overhead. Furthermore, theoretical analysis of privacy-utility trade-offs in the context of heat kernel-enhanced federated clustering remains absent from the literature.
2.2 Research Objectives and Contributions
Motivated by these gaps, this work pursues the following research objectives:
Objective 1: Develop a theoretically grounded framework that integrates heat kernel coefficients—derived from quantum field theory and spectral graph analysis—into multi-view fuzzy clustering, enabling geometry-aware similarity assessment that captures intrinsic manifold structures.
Objective 2: Extend the centralized heat kernel-enhanced framework to federated settings, incorporating differential privacy mechanisms, secure aggregation protocols, and adaptive view weighting strategies tailored for heterogeneous multi-view data across distributed clients.
Objective 3: Establish rigorous mathematical foundations including convergence guarantees for both centralized and federated algorithms, privacy-utility analysis under differential privacy constraints, and communication complexity characterization for federated implementations.
Objective 4: Validate the proposed framework through comprehensive experimental evaluation on synthetic multi-view datasets with controlled geometric complexity, demonstrating substantial improvements over state-of-the-art baselines in clustering accuracy, communication efficiency, and privacy preservation.
Our proposed FedHK-MVFC framework addresses these objectives by integrating heat kernel-based geometric measures with federated multi-view clustering, enabling effective collaboration while preserving data privacy and capturing complex geometric relationships. The framework provides a unified theoretical foundation that bridges quantum field theory concepts with practical federated learning implementations, offering both algorithmic innovations and rigorous mathematical analysis that advance the state-of-the-art in privacy-preserving distributed multi-view clustering.